OpenStudy (anonymous):

If the heat of combustion for a specific compound is -1320.0 kJ/mol and its molar mass is 71.83 g/mol, how many grams of this compound must you burn to release 543.00 kJ of heat?

OpenStudy (anonymous):

Use dimensional analysis. You want your answer in grams, you have the heat released in kJ. So you need to set it up something like\[543=xg * \frac{1}{71.83}\frac{mol}{g}*1320\frac{ kJ}{mol}\]Then solve for x.